### v3

# On misconceptions about the Brier score in binary prediction models

Abstract.
The Brier score is a widely used metric evaluating overall performance of predictions for binary outcome probabilities in clinical research. However, its interpretation can be complex, as it does not align with commonly taught concepts in medical statistics. Consequently, the Brier score is often misinterpreted, sometimes to a significant extent, a fact that has not been adequately addressed in the literature. This commentary aims to explore prevalent misconceptions surrounding the Brier score and elucidate the reasons these interpretations are incorrect.
1. Introduction: What is the Brier score?
The Brier score [8] is a widely used metric evaluating the accuracy of probabilistic predictions in binary outcomes for clinical research [34, 22]. It assesses the overall performance of prediction models that estimate the likelihood of medical outcomes like disease progression or treatment response.
Given a set of probabilistic predictions $p_{i}$ and corresponding observed outcomes $y_{i}$, the Brier score is defined as:
(1.1) $BS(p,y)=\frac{1}{n}\sum_{i=1}^{n}(p_{i}-y_{i})^{2}.$
where:
  • $n$ is the total number of predictions and observations,
  • $p_{i}$ represents the predicted probability of an event occurring for the $i$-th case (e.g., the probability of a patient developing a condition),
  • $y_{i}$ is the actual observed outcome (coded as 1 if the event occurred, 0 if it did not).
Typically, the $y_{i}$ in such prediction models are assumed to be realisations of independent Bernoulli random variables $Y_{i}\sim Bern(q_{i})$, where $q_{i}\in[0,1]$ [22, 33]. Simplifying, one can think of a Bernoulli variable like a potentially biased coin flip: each patient either has the event 1 or 0, based on their individual risk. Two random variables are independent if the probability of an outcome for one is unaffected by the outcome of the other, for a formal explanation see, e.g., [41]. Correspondingly the best i-th prediction that can be obtained is the true underlying probability, i.e, $p_{i}=q_{i}$.
Brier score offers a comprehensive evaluation of probabilistic predictions and is a strictly proper scoring rule [9, Theorem 1], meaning that it incentivizes honest probabilistic forecasting as in expectation it is minimised if and only if the predictions are the true probabilities [20]. Note that there is a difference between good predictions with respect to true probabilities versus usefulness of the predictions in a clinical context: While a clinician might find it easiest to work with a classification of patients into zeros and ones or something very close to that, the quality of a probabilistic prediction model is judged based on how close the predictions are to the true probabilities. As the Brier score is strictly proper, the expectation of Brier score is minimised if and only if the predictions correspond to the true probabilities. If desired, based on the predictions, as a second step the predictions can be used to derive and evaluate a classification [21] or analysed with respect clinical consequences, e.g., via net benefit [36]. However, the qualities such evaluations judge is different to the Brier score [5, 21], in particular these are not strictly proper.
Despite its widespread use [37, 28, 31], the Brier score is potentially often misunderstood in clinical research. Unlike more familiar statistical notions, it does not fit neatly into traditional statistical concepts potentially commonly taught in medical education [25, 4, 30, 32]. Furthermore, the Brier score is mathematically equivalent to the mean squared error (MSE), a concept introduced by C.F. Gauss [18]. The MSE is widely applied in ordinary least squares regression [12, §$\S$ 11.3.1], statistical learning [26, §$\S$ 2.2.1], or machine learning evaluation [15]. The connection of Brier score and MSE can also lead to confusion. Brier score and MSE are used in different contexts, as the Brier score compares a probability to an outcome of the binary random variable in the sense of scoring rules [20], while the MSE usually compares two real continuous values, in statistics typically comparing an estimator to the true value [12]. In particular, misconceptions about the Brier score are not uncommon and can sometimes be reinforced by potentially misleading statements in the literature [35, 36, 34, 1, 40, 11]. Given the importance of accurate interpretation in clinical applications, it is crucial to address these misunderstandings.
Several works have explored the evaluation of binary predictive models. [36] thoroughly reviews recent and more traditional measures for prediction models, Harrell [22, §$\S$ 10] provides a comprehensive overview of regression model performance metrics, Steyerberg [34] offers far-reaching guide to prediction model building in medicine, while [42] discusses the Gini coefficient, Pietra index, and other classification measures. Other studies, such as [24], examine alternative metrics designed to improve model performance assessment, and [6] compares the Brier score with net benefit analysis in diagnostic testing. Furthermore there are various decompositions of the Brier score [43], some can also be used to construct statistical tests for the Brier score [33]. Brier score also remains significant in AI-driven medical predictions [10], and can also be used for evaluation of survival outcomes [1].
The evaluation of binary prediction models has been extensively addressed in the medical literature. Steyerberg et al.[36] provide a comprehensive review of traditional and more recent performance measures for clinical prediction models. Harrell offers a detailed discussion of regression model evaluation metrics [22]. Steyerberg also presents a thorough guide to the development and validation of clinical prediction models [34]. Wu and Lee discuss alternative metrics such as the Gini coefficient and Pietra index [42]. Other studies have focused on expanding the toolbox for performance assessment, including alternative metrics tailored to clinical utility [24]. Assel et al. compared the Brier score with decision-analytic approaches like net benefit analysis in diagnostic test evaluation [6].
Further, various decompositions of the Brier score have been proposed to better understand its components [43, 33]. The Brier score continues to be an important metric in AI-based predictive modeling in healthcare[10], and applies equally to the evaluation of prognostic models for survival outcomes [1].
This commentary aims to clarify common misunderstandings about the Brier score, explain why they arise, and provide guidance on its appropriate interpretation in clinical research. As it is also directed at researchers in medicine, we first introduce some notions, after which we give a summary of the misconceptions.
Key terms we will use and what they mean
  • Random variable: A random variable is a quantity that depends on the outcome of a random process. For example, let $Y$ denote whether a leaving patient is readmitted within 30 days:
    $Y=\begin{cases}1&\text{if readmitted}\\ 0&\text{if not readmitted}\end{cases}$
    Before observation, $Y$ is unknown and varies due to chance. If $Y\sim Bern(q)$ for $q=0.2$, the probability to observe a $1$ is $20\%$, and the probability to observe a 00 is $80\%$.
  • Expectation: Expectation is a way to describe the average outcome we expect over the long run, if we repeat a situation many times under the same conditions, denoted by $\mathbb{E}(\cdot)$. The expectation of $Y\sim Bern(q)$ for $q=0.2$, $\mathbb{E}(Y)$, is $0.2$.
  • Perfect prediction: In clinical prediction models, a perfect prediction means that the predicted probability of an outcome exactly matches the true underlying risk for each individual patient. In the case of our setting of the Brier score, if the true probabilities are given by $q=(q_{1},\cdots,q_{n})$, the perfect prediction is given by $p=q$.
    Suppose a model estimates that a patient has a 20% chance of being readmitted to the hospital within 30 days. If that patient’s true risk is also exactly 20%, then the model has made a perfect prediction for that individual.
Quick reference for practitioners for the misconceptions
Misconception Reality
# 1: Brier score of 0 = perfect model A Brier score of 0 implies extreme predictions (0% or 100%) that exactly match outcomes. This is odd in practice and may indicate overfitting or an error.
# 2: Lower Brier score always means a better model A lower Brier score can be misleading across datasets with different underlying distributions for the outcomes. It is only meaningful to compare Brier scores within the same population and context.
# 3: A low Brier score indicates good calibration Calibration and Brier score measure different aspects. Calibration refers to how well predicted probabilities reflect observed risks; a model can have a low Brier score and still be poorly calibrated.
# 4: A Brier score near $\bar{y}-\bar{y}^{2}$ means the model is useless Even perfect predictions can yield a Brier score near $\bar{y}-\bar{y}^{2}$ if the true risks are close to the mean incidence. This does not imply non-informativeness.
# 5: Brier score cannot exceed $\bar{y}-\bar{y}^{2}$ for reasonable predictions This is a common belief, but in small samples and some underlying distribution, the Brier score can exceed the threshold due to chance or reasonable but not perfect predictions.
Table 1. Common misconceptions about the Brier score and how they contrast with statistical reality.
Related literature
Some of our points have been previously observed. We review related references that observe similar findings. However, given the widespread use of the Brier score, our literature review is necessarily partial. [27] outlines examples where the model comparison of expected BSs is potentially contrary to how a human would judge. [33] outlined the distinction between calibration and prediction errorand the misunderstanding that lower Brier score indicates better calibration.
Acknowledgements
We thank Lucia de Andres Bragado, Julien Vionnet, Simon Schwab, and Matthew Parry for helpful discussions and feedback.
2. More on the BS
In this section, we review key properties of the BS, delve into some mathematical points, and discuss the means we will use to understand the BS.
As a first point, when considering the Brier score of (1.1), there are two simple related BSs. On the one hand, one can compare (1.1) to the Brier score when entering $p_{1/2}=(1/2,\cdots,1/2)$, giving $BS(p_{1/2},y)=1/4$. On the other hand, we can compare (1.1) to the Brier score when entering the mean incidence $\bar{y}_{v}=(\bar{y},\cdots,\bar{y})$, which gives
$BS(\bar{y}_{v},y)=\bar{y}-\bar{y}^{2}.$
Clearly, $BS(p_{1/2},y)=1/4\geq BS(\bar{y}_{v},y)=\bar{y}-\bar{y}^{2}$. Both for very high or very low incidences, $BS(\bar{y}_{v},y)$ is low. Furthermore we note that $\bar{y}-\bar{y}^{2}$ si symmetric around $0.5$, see Figure 1, and e.g. for $\bar{y}=0.1$ or $\bar{y}=0.9$, $BS(\bar{y}_{v},y)=0.09$.
2.1. Key properties about the BS
We summarise essential insights into the Brier score.
  1. (BI)
    Range and interpretation: The Brier score takes values in the interval $[0,1]$, with lower values typically indicating more accurate probabilistic predictions.
  2. (BII)
    The Brier score is a random variable: The Brier score is a function of random variables and as such it is also a random variable. In clinical prediction models, the outcomes $y_{i}$ are realisations of $Y_{i}\sim\text{Bern}(q_{i})$, where $q_{i}\in[0,1]$.
  3. (BIII)
    Optimal predictions and true probabilities: The (unique) optimal predictions minimising the expected Brier score are the true outcome probabilities, i.e., $p_{i}=q_{i}$, as the Brier score is a strictly proper scoring rule [9, Theorem 1].
  4. (BIV)
    Expectation of Brier score for $n=1$: Knowing the true probability, we can calculate the expectation. For $n=1$, let $Y_{1}\sim Bern(q_{1}),q_{1}\in[0,1]$ with prediction $p_{1}\in[0,1]$. Then the expectation of Brier score is given as
    (2.1) $g(p_{1},q_{1}):=\mathbb{E}_{Y_{1}\sim Bern(q_{1})}[BS(p_{1},Y_{1})]=p_{1}^{2}-2p_{1}q_{1}+q_{1}.$
    For the interested reader, the calculation is given in Appendix B.1. The optimal prediction minimising (2.1) is $p_{1}=q_{1}$. We next go through two key cases
    • For perfect prediction, the expectation value is given by
      (2.2) $f(q_{1})=q_{1}-q_{1}^{2}.$
      As examples, consider the following cases:
      • if $q_{1}=1/2=p_{1}$, the expected Brier score is $f(1/2)=1/4$ corresponding to the maximum of (2.2),
      • if for $q_{1}=1/10=p_{1}$, $f(1/10)=9/100$.
      The expectation of the Brier score as a function of $q_{1}$, when $p_{1}=q_{1}$, i.e., (2.2), looks as follows:
      $p$$f(p)$10.50.25
      Figure 1. Expectation of the Brier score for the optimal prediction as a function of the underlying true probability with $p_{1}=q_{1}$.
    • Next we compare the expected Brier score under perfect prediction with $p_{1}=q_{1}$ to
      • the expectation of the Brier score under perfect prediction but increased true probability, i.e., $\tilde{p}_{1}:=p_{1}+\varepsilon=\tilde{q}_{1}=q_{1}+\varepsilon$, and taking the difference
        (2.3) $g(q_{1}+\varepsilon,q_{1}+\varepsilon)-g(q_{1},q_{1})=\varepsilon(1-2q_{1}-\varepsilon)$
        Note that the difference depends strongly on $q_{1}$.
      • to the expectation of the Brier score with same true probability but slightly wrong prediction $\tilde{p}_{1}:=p_{1}+\varepsilon,\tilde{q}_{1}=q_{1}\in[0,1]$ and taking the difference
        (2.4) $g(q_{1}+\varepsilon,q_{1})-g(q_{1},q_{1})=\varepsilon^{2}$
        Note that the difference depends does not depend on $q_{1}$, and the same holds for $\tilde{p}_{1}:=p_{1}-\varepsilon$. The calculations are in Appendix §$\S$ B.2.
    • From the previous point, we conclude that if we slightly predict wrong, the costs are almost inexistent, e.g. if the true probability is $q_{1}$ but we predicted $p_{1}=q_{1}+0.1$, the expected difference between the Brier score of the perfect prediction and the slightly wrong one is $0.01$ by (2.4). However, if the true probability is changed e.g. from $q_{1}=0.1$ to $\tilde{q}_{1}=0.2$ and we have perfect prediction, the expected difference of the BSs is $0.07$ by (2.3).
    • Furthermore, we note the following nice order preservation property from (2.4): If $p_{1},\tilde{p_{1}}\in[0,1]$ such that $|p_{1}-q_{1}|<|\tilde{p_{1}}-q_{1}|$, then
      $\mathbb{E}_{Y_{1}\sim Bern(q_{1})}[BS(p_{1},Y_{1})]<\mathbb{E}_{Y_{1}\sim Bern(q_{1})}[BS(\tilde{p_{1}},Y_{1})].$
      This property is known as effective in the literature on scoring rules, cf., e.g., [17].
  5. (BV)
    Dependence of expectation and distribution of Brier score on true probabilities: The Brier score with several observations as in (1.1) consists of the mean value of the one-dimensional BS’s. The expectation of Brier score in (1.1) and its distribution depend on the true outcome probabilities $q_{i}$, which are generally unknown, even when these are used as predictions. Based on (2.2) or Figure 1 its follows e.g. that when the $q_{i}$ are mostly concentrated around zero and one, then the expectation of the Brier score under perfect predictions (1.1) is close to zero. On the other hand, when they are mostly concentrated around e.g. $0.5$, the expectation of the Brier score (1.1) with perfect predictions is close to $0.25$.
  6. (BVI)
    Unobservability of true probabilities in clinical data: In practice, the true probability of an event occurring for an individual patient is not directly observable. Each patient is unique, and we only observe whether the event occurs or not (i.e., a binary outcome). Note however, that in simulation settings one can compare the true probabilities to estimated probabilities of regression or ML models, i.e. as done here [39, 13].
In particular we note that although individual-level true probabilities will remain unknown, we can approximate their average value across a population by calculating the mean of observed outcomes in overall, or in similar patient groups. This provides a practical way to assess model calibration in clinical settings via calibration in the large (CIL) [7, 38]
(2.5) $CIL(p,y)=\frac{1}{n}\sum_{i=1}^{n}p_{i}-\frac{1}{n}\sum_{i=1}^{n}y_{i}.$
2.2. Notable mathematical features about the BS
Below, we summarise basic mathematical properties of the Brier score for the understanding of the reader. Mathematical details and arguments are given in Appendix B.3.
Consider the Brier score of $n$ outcomes, where each $Y_{i}\sim Bern(q_{i})$. Let the expected incidence be denoted by $\bar{q}$, i.e. $\bar{q}:=\frac{\sum_{i=1}^{n}q_{i}}{n}$.
  1. (MI)
    Brier score of non-informative model that uses prevalence as prediction, i.e. $p=(\bar{y},\cdots,\bar{y})$: The Brier score of (1.1) with the non-informative mean as predictor $\bar{y}_{v}=(\bar{y},\cdots,\bar{y})$, is given by $\bar{y}-\bar{y}^{2}$, i.e.
    $BS(\bar{y}_{v},y)=\bar{y}-\bar{y}^{2}.$
  2. (MII)
    Bound on expectation of perfect prediction, i.e. for $p=(q_{1},\cdots,q_{n})$: The expected value of the average Brier score of (1.1) with the true probabilities as predictors, i.e., $p=(q_{1},\cdots,q_{n})$, is bounded above by $\bar{q}-\bar{q}^{2}$, i.e.,
    $\mathbb{E}[BS((q_{1},\cdots,q_{n}),(Y_{1},\cdots,Y_{n}))]\leq\bar{q}-\bar{q}^{2},$
    with equality if and only if $q_{i}=\bar{q}$ for all $i\in[n]$.
  3. (MIII)
    Typical Brier score with perfect prediction for large $n$: Assume the true probabilities $q_{i}$ itself are realisations of random variables $Q_{i}\sim F$. Then, by the law of large numbers(LLN) the Brier score (1.1) for $n$ big roughly equals $\mathbb{E}[BS(Q_{1},Y_{1})]$, and probabilities for deviation from this value can be calculated via the central limit theorem (CLT). Hence $\mathbb{E}[BS(Q_{1},Y_{1})]$ roughly equals $BS((q_{1},\cdots,q_{n}),(y_{1},\cdots,y_{n}))]$ for $n$ large, and $\bar{y}$ roughly equals $\bar{q}$. More detail on this perspective is in Appendix §$\S$ A.
2.3. Main takeaways of the observations on the BS
To summarise our previous points, we conclude that an observed Brier score is a function of
  1. (I)
    the underlying true probabilities (the $q_{i}$s),
  2. (II)
    the closeness of the predictions when compared to the true probabilities (how close $p_{i}$ is from $q_{i}$),
  3. (III)
    some randomness that comes from the Bernoulli random variables (the observed $y_{i}$ that are realisations of $Y_{i}\sim Bern(q_{i})$).
The influence of the randomness decreases with $n$ (in idealised proper settings by LLN and CLT), and the expectation of the Brier score can be seen as the long term average, which represent typical values if $n$ is big.
2.4. Means of understanding the BS
We will use the following approaches to better understand the Brier score (1.1):
  • Expectation of the BS: We will analyze the expected value the Brier score (1.1) takes, which will help to illustrate typical observed values of the BS.
  • Simulation-Based Evaluations: We assess the BS’s behavior by simulating Bernoulli outcomes with different probability distributions for $q_{i}$ and different sample sizes. This approach intuitively explores properties in both realistic and idealized medical prediction scenarios.We will use the following elements in our simulations, where we give more detail in the Appendix §$\S$ E via the Aims, Data-generating mechanism, Estimands and other targets, Methods, Performance measures (ADEMP) framework [29].
    • Sample size $n$:. Settings considered: $n\in\{300,1000\}$.
    • True distribution: The $q_{i}\in[0,1]$ used to simulate $Y_{i}\sim Bern(q_{i})$. The $q_{i}$ itself are considered as realisations of random variables.
    • Predictor distribution: The $p_{i}\in[0,1]$ used in (1.1). These are considered as functions of $q_{i}$ with potentially random error.
    • Estimand: Includes median Brier score and $2.5\%$ resp. $97.5\%$ quantiles.
3. Misconceptions
Below are the most common misinterpretations of the Brier score when evaluating probability predictions for binary events, accompanied by an explanation and examples illustrating why it is incorrect.
3.1. Misconception #1: A Brier score of 0 means a perfect model, and a perfect model has Brier score 0.
  • A Brier score of 0 means a perfect model. A Brier score of 0 implies perfect alignment between predicted probabilities and observed outcomes. While the true probabilities are unknown, they lie within $[0,1]$, and we usually expect that a majority is in $(0,1)$. Hence, rather than signalling a perfect model, an observed Brier score of 0 potentially indicates a serious error. Given the binary outcomes, this would imply predictions were exclusively 0 or 1, contradicting the goal of probability estimation.
  • A perfect model has Brier score 0. Note that under the assumption that at least one of the true probabilities is in $(0,1)$, a Brier score of 0 is mathematically impossible perfect predictions (see Appendix C). Consider expected values of a perfect model. If $n=1$, and say $q_{1}=0.1$, this expectation is not equal to zero for the perfect model by (2.2), nor can an observed Brier score be equal to zero. If we evaluate several perfect predictions, the corresponding expected value of the Brier score is the mean over (2.2) which is not zero (unless for all $i$, $q_{i}\in\{0,1\}$).
To illustrate this and for later reference, we present simulations with perfect predictions, showing that even ideal models do not yield a Brier score of 0. In fact, expected Brier scores for a perfect model can be notably high.
True Predictor n Median of BS
distribution distribution (2,5%-&97,5%- Q)
Unif(0,1) Perfect 300 $0.167(0.144,0.19)$
Unif(0,1) Perfect 1000 $0.167(0.154,0.179)$
Unif(0,1) $+Unif(-0.2,0.2)$ 300 $0.179(0.156,0.203)$
Unif(0,1) $+Unif(-0.2,0.2)$ 1000 $0.179(0.165,0.193)$
Smoke prediction Perfect 300 $0.168(0.145,0.191)$
Smoke prediction Perfect 1000 $0.168(0.156,0.181)$
Smoke prediction Biased +0.1 300 $0.178(0.161,0.196)$
Smoke prediction Biased +0.1 1000 $0.178(0.169,0.188)$
Osteo prediction Perfect 300 $0.054(0.037,0.074)$
Osteo prediction Perfect 1000 $0.054(0.045,0.065)$
Osteo prediction Biased +0.1 300 $0.064(0.05,0.08)$
Osteo prediction Biased +0.1 1000 $0.064(0.057,0.073)$
Beta(2,2) Perfect 300 $0.2(0.18,0.22)$
Beta(2,2) Perfect 1000 $0.2(0.189,0.211)$
Table 2. Comparison median of Brier scores under different simulation settings
3.2. Misconception #2: When comparing two prediction models, the model with lower Brier score is better.
Prediction models can be compared across datasets through Brier score, but caution is essential. While a lower Brier score for one prediction model compared to another may suggest better predictive performance, this can be deceptive. Below, we present three scenarios with decreasing levels of caution, each illustrating why direct comparison of BSs may be misleading.
3.2.1. Comparing BSs Across Different Datasets with different incidences:
If the true outcome distributions differ, even perfect models will yield different Brier score distributions and expectations. Thus, cross-dataset comparisons may lack meaningful insights. Observing different incidences indicates that the true outcome distributions differ, making the Brier score comparison unreliable. The Brier score is a function of the true underlying probabilities, the predictions and some randomness, with the true probabilities having a big influence. Hence a lower Brier score in prediction model $A$ when compared to prediction model $B$ does not necessarily indicate a superior model in $A$ with more accurate predictions, as the expectation of Brier score depends on the true probabilities (BV). Compare, e.g. the osteoporosis prediction with $0.1$ bias to the perfect smoke prediction model from Table 5. The biased osteoporosis prediction has median Brier score $0.064$, while perfect smoke prediction has median Brier score $0.168$, nonetheless the perfect model is obviously better.
3.2.2. Comparing BSs Across Different Datasets with very similar or same incidences (also known as class imbalance) is meaningful.
Some recommend to compare Brier score of two prediction models on datasets with similar class imbalances [11]. However, the true probability distribution is unobservable, and by (BV) the true values $q_{i}$ determine the expectation of the Brier score. Thus, even with identical class imbalances, we cannot assume that the true probability distributions are comparable and the same caution as in the previous point should be exercised in interpreting such comparison. Descriptions in population characteristics can indicate similarities and differences of true probabilities. Compare the example with underlying distribution $Unif(0,1)$ where the prediction has error $+Unif(-0.2,0.2)$ to the the example with $Beta(5,5)$ under perfect predictions from Table 5. Both $Unif(0,1)$ and $Beta(5,5)$ give expected incidences of $50\%$. The Brier score of predictions with errors but underlying distribution $Unif(0,1)$ has median Brier score $0.179$, while perfect prediction for underlying $Beta(5,5)$ has median Brier score $0.200$, nonetheless the perfect model is obviously better.
3.2.3. Comparing Models on the same dataset via Brier score:
Even when evaluating two models on the same dataset, a lower Brier score in one does not always strictly imply that this model is better. There are two potential issues:
  1. (I)
    It is possible that the model that is worse has a better score by pure chance. However, with higher sample sizes this becomes more unlikely by the CLT (MIII).
  2. (II)
    The way in which the Brier score in expectation judges which model is better might not align with how a human might judge. In expectation, a change of the prediction from the perfect prediction $q_{i}$ to $q_{i}+\varepsilon$ or $q_{i}-\varepsilon$ gives the same result by (2.4). However, for humans the direction can matter as well, particularly for $q_{i}$ close to zero or one, as observed in [27]. Consider the example with one observation from [27, §$\S$ 3] where the probability of an event is $P(Y_{1}=1)=\frac{1}{10}$ (i.e., 10% chance), and compare two models:
    • Model 1: $p_{1}=0$ (predicting the event will never occur)
    • Model 2: $\tilde{p_{1}}=\frac{1}{4}$ (predicting the event occurs with a probability of 25%)
    Using (2.1), model 1 has an expected Brier score of $0.1$, as compared to $0.1125$ for model 2. As model 1 predicts a zero probability for the observation that has actually a 10% probability, at least intuitively, 25% might seem better for human judgement.
3.3. Misconception #3: A low Brier score indicates good calibration.
A low Brier score does not necessarily indicate good calibration of a model. As we have repeated, the Brier score is a function of the true underlying probabilities, the predictions and some randomness, with the true probabilities having a big influence. Hence it is possible that we have perfect predictions, but low or high Brier score due to the underlying probabilities, or similary not so accurate predictions but a Brier score that is low or high Brier score due to the underlying probabilities, see (BIV). As in expectation a change of the prediction from the perfect prediction $q_{i}$ to $q_{i}+\varepsilon$ or $q_{i}-\varepsilon$ gives the same result by (2.4), miscalibration where errors tend to go mostly in one direction are not punished more. Hence in the examples below from Table 3 we observe that median Brier score from the prediction biased by $0.1$ are the same as the one from the prediction with $+(1-2Bern(1/2))\cdot 0.1$ which has prediction errors that are as often and as much higher as they are lower than the true probability..
Assessing calibration should be done using additional metrics like CIL, calibration curves or similar evaluation components [7, 38]. Therefore, while the Brier score is a useful measure, it should not be the sole criterion for evaluating model performance.
True Predictor n Median CIL Median Brier score
distribution distribution (2,5%-&97,5%- Q) (2,5%-&97,5%- Q)
Unif(0,1) Perfect 1000 $0(-0.025,0.025)$ $0.166(0.154,0.179)$
Unif(0,1) $+0.1$ 1000 $0.095(0.07,0.12)$ $0.176(0.163,0.189)$
Unif(0,1) $+(1-2Bern(1/2))\cdot 0.1$ 1000 $0(-0.026,0.026)$ $0.176(0.163,0.189)$
Osteo model Perfect 1000 $0(-0.015,0.014)$ 0.054 (0.045,0.065)
Osteo model $+0.1$ 1000 $0.1(0.085,0.115)$ $0.064(0.057,0.073)$
Osteo model $+(1-2Bern(1/2))\cdot 0.1$ 1000 $0.03(0.015,0.045)$ $0.061(0.051,0.071)$
Table 3. Comparison of expectation of Brier score and quantiles under different simulation settings (Monte-Carlo simulation with $5000$ runs per setting)
3.4. Misconception #4: Having a Brier score of around $\bar{y}-\bar{y}^{2}$ where $\bar{y}$ is the mean observed incidence means we have a useless or non-informative model. As an example, if $\bar{y}=0.5$, then $\bar{y}-\bar{y}^{2}=0.25$, or , if $\bar{y}=0.1$, then $\bar{y}-\bar{y}^{2}=0.090$
Recall from (BIII) that the lowest expected Brier score occurs when predicted probabilities match the true probabilities, representing a perfect prediction that cannot be improved. However, we cannot observe the true probabilities. In some cases of observed incidences around $0.5$, perfect predictions yield a Brier score of close to 0.25, as shown in the following simulations. Hence if we observe a Brier score of around $\bar{y}-\bar{y}^{2}$, besides having a relatively bad model, the following alternative explanations could explain such a BS:
  • Many of the true probabilities are around $\bar{y}$ or $1-\bar{y}$, which can make the expected Brier score of perfect predictions close to $\bar{y}-\bar{y}^{2}$ (MII).
  • If $n$ is relatively low, as the Brier score is a realization of a random variable (BII), bigger random error are likely which can make the Brier score higher than its expectation.
Only if we can assume that neither of the above is excluded can conclude that the prediction model used is relatively bad.
As an illustration of example values, consider the median $\bar{y}-\bar{y}^{2}-BS_{perf}$ in Table 4, where $BS_{perf}$ is the Brier score under perfect predictions, which depends entirely on the underlying distribution, and again these medians strongly depend on the underyling distribution, but can be very low.
True Predictor n Median of $\bar{y}-\bar{y}^{2}-BS_{perf}$
distribution distribution (2,5%-&97,5%- Q)
Unif(0,1) Perfect 300 $0.083(0.059,0.105)$
Unif(0,1) Perfect 1000 $0.083(0.07,0.095)$
Beta(5,5) Perfect 300 $0.022(0.006,0.037)$
Beta(5,5) Perfect 1000 $0.022(0.014,0.031)$
Osteo model Perfect 300 $0.01(0.001,0.021)$
Osteo model Perfect 1000 $0.011(0.005,0.016)$
Smoke model Perfect 300 $0.026(0.01,0.041)$
Smoke model Perfect 1000 $0.026(0.017,0.034)$
Table 4.
3.5. Misconception #5: For an observed incidence of $\bar{y}$, the Brier score (1.1) for reasonable predictions can not be bigger than $\bar{y}-\bar{y}^{2}$
While a priori we would not expect the Brier score for reasonable predictions to be bigger than $\bar{y}-\bar{y}^{2}$, it can happen. As we can not observe the true probabilities (BVI) we can not exclude the possibility that the expected Brier score for perfect predictions is close to the (unobservable) expected incidence $\bar{q}:=\frac{\sum_{i=1}^{n}q_{i}}{n}$, which is the expectation of $\bar{Y}$. Recall again from (BII) that the Brier score is a random variable. This randomness comes from the outcomes, making also the mean incidence $\bar{y}$ a random variable. All we know is that for perfect predictions, the probability to observe $BS>\bar{y}-\bar{y}^{2}$ will go to zero as $n\to\infty$ by (MIII). However, as in practice we will not be able to derive the perfect predictions, the expectation will be higher and this argument does not apply.
Hence as in the previous point, a Brier score above $\bar{y}-\bar{y}^{2}$ can be obtained when, apart from having a relatively bad model, the following alternative cases apply:
  • Many of true probabilities are around $0.5$, or tend to be close to each other.
  • If $n$ is relatively low, as the Brier score is a realization of a random variable (BII), random error can make the Brier score higher than its expectation by chance.
To illustrate this, we show that even for perfect predictions in some settings with low $n$, the probability of having a Brier score that is bigger than $\bar{y}-\bar{y}^{2}$ can be nonzero. As in practice we will not have perfect predictions, in the settings below for reasonable predictions the probability to have a Brier score bigger than $\bar{y}-\bar{y}^{2}$ is bigger, and the estimations can be understood as lower bounds for corresponding probabilities.
True Predictor n Event fraction
Distribution Distribution where $\bar{y}-\bar{y}^{2}>BS$
Unif(0,0.2) Perfect 300 0.060
Unif(0.3,0.7) Perfect 300 0.030
Osteroporosis model Perfect 300 0.019
Table 5. Comparison of expectation of Brier score and quantiles under different simulation settings (Monte-Carlo simulation with $5000$ runs per setting)
4. Connections to some other scores
We mention three other scores connected to BS. The Brier score from (1.1) equals mathematically the MSE, hence its square root is the root mean squared error (RMSE):
(4.1) $RMSE(p,y):=\sqrt{BS(p,y)}.$
As the square root on $[0,1]$ is order preserving, i.e., if $a,b\in[0,1],a\leq b$ then $\sqrt{a}\leq\sqrt{b}$, and bijective, RMSE also takes values in $[0,1]$ and most of the observations we made apply sililarly to RMSE.
Two similar and often-used scores are the mean-absolute error (MAE) that is defined as
(4.2) $MAE(p,y)=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-y_{i}|,$
and the CIL [34]
(4.3) $CIL(p,y)=\frac{1}{n}\sum_{i=1}^{n}p_{i}-y_{i}.$
These relate to Brier score (or RMSE) through the following inequalities [3]
(4.4) $CIL(p,y)\leq MAE(p,y)\leq RMSE(p,y)$
as well as
(4.5) $MSE(p,y)\leq MAE(p,y)\leq RMSE(p,y).$
In particular, we note that both MAE and CIL are not strictly proper, which we also outline in the Appedix D.
5. Conclusions and final remarks
We adressed common misconceptions regarding the interpretation and use of the Brier score. Summarising, the observed Brier score is a function of
  1. (I)
    the underlying true probabilities,
  2. (II)
    the closeness of the predictions when compared toe the true probabilities,
  3. (III)
    some randomness that comes from the Bernoulli random variables.
In particular, the effect of the true underlying probabilities on the expectation of Brier score is strong, potentially stronger than the closeness of the predictions to the true probabilities as outlined in §$\S$ 2.1. We showed that a Brier score of zero is rather an indication of an error in realistic settings, and a very low Brier score does not necessarily indicate a perfect model. Indeed, the Brier score should be interpreted with caution, having in mind the relation to the true probabilities and the characteristics of the dataset. Direct comparisons of BSs across models on the same data can be done, and across different data should be avoided, or if done interpreted carefully. Additionally, we stressed that a low Brier score does not guarantee good calibration and should be complemented with additional calibration metrics. Furthermore the randomness in the Brier score, through the CLT, decreases as the sample size increases. A recent literature review on clinical prediction models [14] found that sample sizes used had median sample size $1250$ with $(Q1,Q3)=(353,188860)$. Hence at least a quarter of the prediction models had relatively low sample sizes, with randomness similar to the setting in our simulations with $n=300$.
Overall, while the Brier score remains a valuable metric for assessing probabilistic predictions, its interpretation requires an understanding of how the underlying probabilities and closeness of predictions influence the observed BS. Once misconceptions are avoided, the Brier score serves as a reliable relative measure of overall performance. In particular, it is strictly proper and hence in expectation, it is minimised uniquely at the true probabilities, making relative comparisons on the same data meaningful [9, Theorem 1], besides the nice order preservation property from (BIV).
Appendix A Why we care about the expectation: law of large numbers and central limit theorem
Another way to understand the behavior of the Brier score for perfect predictions is through basic mathematical tools. Consider the setting of the simulations considered, where the true probability $q_{i}$ itself is a realisation of a random variable $Q_{i}\sim F$. Denote by $J_{n}=(Q_{1},\cdots,Q_{n})$ the random vector of the first $n$ true probabilities and $Z_{n}=(Y_{1},\cdots,Y_{n})$ the realisation of the corresponding $n$ Bernoulli random variables. We can consider the Brier score of the optimal predictor $J_{n}$ as
$BS(J_{n},Z_{n})\xrightarrow[]{n\to\infty}\mathbb{E}[(Q_{1}-Y_{1})^{2}]$
For sufficiently large $n$, the observed Brier score provides a stable estimate of its expectation by the LLN [19].
Furthermore, the individual terms $(Q_{i}-Y_{i})^{2}$ have finite variance and hence by the CLT the distribution of the BS, when properly normalised, approaches a normal distribution.
[…]
Appendix B More detail for mathematical understanding BS
B.1. Expectation of Brier score in one dimension
Consider $Y_{1}\sim Bern(q_{1}),q_{1}\in[0,1]$ and one prediction $p_{1}\in[0,1]$, then the expected Brier score is given as
$\mathbb{E}[BS(p_{1},Y)]=p_{1}^{2}-2p_{1}q_{1}+q_{1}.$
In order to derive the above formula, we can proceed as follows.The expected Brier score is
$\mathbb{E}[BS(p_{1},Y_{1})]=\mathbb{E}[(p_{1}-Y_{1})^{2}].$
We can expand the terms to get
$\mathbb{E}[(p_{1}-Y_{1})^{2}]=\mathbb{E}[p_{1}^{2}-2p_{1}Y_{1}+Y_{1}^{2}].$
Since $p_{1},p_{1}^{2}$ are constants, we get:
$\mathbb{E}[p_{1}^{2}-2p_{1}Y_{1}+Y_{1}^{2}]=p_{1}^{2}-2p_{1}\mathbb{E}[Y_{1}]+\mathbb{E}[Y_{1}^{2}].$
For a Bernoulli-distributed variable $Y_{1}$:
$\mathbb{E}[Y_{1}]=q_{1},\quad\mathbb{E}[Y_{1}^{2}]=\mathbb{E}[Y_{1}]=q_{1}.$
Hence the expected Brier score is:
$\mathbb{E}[BS(p_{1},Y_{1})]=p_{1}^{2}-2p_{1}q_{1}+q_{1}.$
B.2. Differences in expectation of Brier score in one dimension
Recall that
$g(p_{1},q_{1})=p_{1}^{2}-2p_{1}q_{1}+q_{1}.$
  • To derive the first result, we compute:
    $g(q_{1}+\varepsilon,q_{1}+\varepsilon)-g(q_{1},q_{1})=q_{1}^{2}+2q_{1}\varepsilon+\varepsilon^{2}-2(q_{1}^{2}+2q_{1}\varepsilon+\varepsilon^{2})+q_{1}+\varepsilon-q_{1}+q_{1}^{2}$
    Thus,
    $g(q_{1},q_{1})-g(q_{1}+\varepsilon,q_{1}+\varepsilon)=\varepsilon(1-2q_{1}-\varepsilon)$
  • For the second expression,
    $g(q_{1}+\varepsilon,q_{1})-g(q_{1},q_{1})=-q_{1}^{2}-2q_{1}\varepsilon-\varepsilon^{2}+2q_{1}^{2}+2q_{1}\varepsilon-q_{1}$
    $=q_{1}^{2}-\varepsilon^{2}-q_{1}$
    Thus, the difference:
    $g(q_{1}+\varepsilon,q_{1})-g(q_{1},q_{1})=\varepsilon^{2}$
B.3. Understanding the expectation of the general BS
In case $p_{1}=q_{1}$, the expectation value of Brier score is given by
(B.1) $f(p_{1})=p_{1}-p_{1}^{2}.$
This is a strictly concave function, meaning that for any $\alpha\in[0,1]$ and any $x,y\in[0,1]$,
(B.2) $f((1-\alpha)x+\alpha y)\geq(1-\alpha)f(x)+\alpha f(y)$
Now we come back to the case where we have $n$ observations and we consider the BS.
Lemma B.1.
Consider the Brier score of $n$ outcomes, where $\frac{\sum_{i=1}^{n}y_{i}}{n}=\bar{y}$. Then the Brier score of (1.1) with prevalence as predictors, i.e., $(p_{1},\cdots,p_{n})=(\bar{y},\cdots,\bar{y})$ equals $\bar{y}-\bar{y}^{2}$, i.e.
$BS((\bar{y},\cdots,\bar{y}),(y_{1},\cdots,y_{n}))=\bar{y}-\bar{y}^{2}$
Proof.
Let $n=a+b$ such that $\bar{y}=\frac{a}{a+b}$, and Brier score is given as
$\frac{1}{a+b}(a(\frac{a}{a+b}-1)^{2}+b(\frac{a}{a+b})^{2})$
which we rewrite as
$\frac{a}{a+b}(\frac{a}{a+b}-1)^{2}-(\frac{a}{a+b}-1)(\frac{a}{a+b})^{2})=\bar{y}(\bar{y}-1)^{2}+(1-\bar{y})\bar{y}^{2}=\bar{y}-\bar{y}^{2}$
Lemma B.2.
Consider the Brier score of $n$ outcomes, where $\frac{\sum_{i=1}^{n}q_{i}}{n}=\bar{q}$. Then the expected value of the average Brier score of (1.1) with the true probabilities as predictors, i.e., $(p_{1},\cdots,p_{n})=(q_{1},\cdots,q_{n})$, is bounded above by $\bar{q}-\bar{q}^{2}$, i.e.
$\mathbb{E}[BS((q_{1},\cdots,q_{n}),(Y_{1},\cdots,Y_{n}))]\leq\bar{q}-\bar{q}^{2},$
with equality if and only if $q_{i}=\bar{q}$ for all $i\in[n]$.
Proof.
Let $n\geq 1$. Then
$\mathbb{E}[BS((q_{1},\cdots,q_{n}),(Y_{1},\cdots,Y_{n}))]=\frac{1}{n}(\sum_{i=1}^{n}q_{i}-q_{i}^{2}).$
We can rewrite this as
$=\frac{1}{n}\sum_{i=1}^{n}q_{i}-\frac{1}{n}\sum_{i=1}^{n}q_{i}^{2}=\bar{q}-\frac{1}{n}\sum_{i=1}^{n}q_{i}^{2}.$
As the function $x\to x^{2}$ is strictly convex, we can apply Jensens Inequality to get
$\frac{1}{n}\sum_{i=1}^{n}q_{i}^{2}\geq(\frac{1}{n}\sum_{i=1}^{n}q_{i})^{2}=\bar{q}^{2},$
such that finally we can bound it as
$\bar{q}-\frac{1}{n}\sum_{i=1}^{n}q_{i}^{2}\geq\bar{q}-\bar{q}^{2}.$
The statement with equality if and only if $q_{i}=\bar{q}$ for all $i\in[n]$ also follows from Jensen. ∎
Another simple observation, with proof here is the following
Lemma B.3.
Consider the Brier score of $n$ outcomes, where $\frac{\sum_{i=1}^{n}q_{i}}{n}=c$. Then the expected value of the average Brier score of (1.1) with the non-informative mean as predictors, i.e., $(p_{1},\cdots,p_{n})=(c,\cdots,c)$, is given by $c-c^{2}$, i.e.
$\mathbb{E}[BS((c,\cdots,c),(Y_{1},\cdots,Y_{n}))]=c-c^{2}$
Proof.
Using (2.1) and (1.1) we get
$\mathbb{E}[BS((c,\cdots,c),(Y_{1},\cdots,Y_{n}))]=\frac{1}{n}\sum_{i=1}^{n}c^{2}-\frac{1}{n}\sum_{i=1}^{n}2cq_{i}+\frac{1}{n}\sum_{i=1}^{n}q_{i}$
which we can simplify using $\frac{\sum_{i=1}^{n}q_{i}}{n}=c$ to get
$c^{2}-2c\frac{1}{n}\sum_{i=1}^{n}q_{i}+\frac{1}{n}\sum_{i=1}^{n}q_{i}=c^{2}-2c^{2}+c=c-c^{2},$
which is what we wanted to show. ∎
Appendix C Mathematical proof impossibility of Brier score 0
Recall the assumption.
Assumption 1.
Assume at least one of the true probabilities $q_{i}$ are in $(0,1)$.
Lemma C.1.
Let $n\in\mathbb{N}_{\geq 1}$, and let $y=(y_{1},\cdots,y_{n})$ be a realisation of a sequence of independent random variables, where $Y_{i}\sim Bern(q_{i})$. If assumption 1 holds, the Brier score of the perfect prediction $p_{perf}=(q_{1},\cdots,q_{n})$ is bigger than zero, i.e.,
$BS(p_{perf},y)>0.$
Proof.
Denote by $q=(q_{1},\cdots,q_{n})$ the vector of true probabilities, which also determines the perfect prediction vector $p_{perf}=q$. Assume we reordered them such that $q_{1}$ is in $(0,1)$, which holds by assumption assumption 1. Define the following constant
$\varepsilon:=\min\{q_{1},1-q_{1}\}$
By reordering and assumption, $\varepsilon>0$. We can bound $BS(p_{perf},y)$ from below as follows
$0<\frac{\varepsilon^{2}}{n}\leq BS(p_{perf},y).$
Appendix D MAE and CIL are not strictly proper
Recall the formulas from (4.2) and (4.3), and let $>\sim Bern(q),q\in[0,1]$. It is enough to do the case $n=1$ for MAE and CIL, the case for more obsevations follows analogous. We have to show that the expectation is not uniquely minimised (resp maximised if the orientation is opposite) for $p\in[0,1]$ if we choose $p=q$ [20].
  • We calculate
    $\mathbb{E}[MAE(p,Y)]=qp+(1-q)(1-p),$
    to simplify further let $q=1/2$, then $\mathbb{E}[MAE(p,Y)]=1/2,$ and the expectation does not depend on $p$.
  • We calculate
    $\mathbb{E}[CIL(p,Y)]=p-q$
    which is clear.
Appendix E More details for the simulation via the ADEMP framework
We give more details on the simulations used in ADEMP framework [29]. It took roughly 1h to run it on a Mac Studio M2 Max 2023.
E.1. Aims
  • The main aim is to compare Brier score across different settings both for true probabilities as well as predictions, which are based on the perfect prediction with some noise or bias added through median.
  • A secondary aim is to compare Brier score to the mean incident BS, to compare it to CIL. In particular to estimate the probability that $\bar{y}-\bar{y}^{2}>BS_{perf}$, and to estimate median of $\bar{y}-\bar{y}^{2}-BS_{perf}$.
E.2. Data Generating Mechanism
E.2.1. $y_{i}$ entered in the BS
We sample true values for $q_{i}$ under some distributions which are subsequently used to simulate $Y_{i}\sim Bern(q_{i})$. The sample distribution for $q_{i}$ are based on the following:
  • $Unif(a,b)$, where $(a,b)\in\{(0,1),(0.3,0.7),(0,0.2)\}$.
  • $Beta(\alpha,\beta)$, where $(\alpha,\beta)\in\{(2,2),(5,5)\}$.
  • Osteoporosis: Logistic regression model based on NHANES 2007/2008 [16] data using the nhanesA-package [2] with complete case analysis; $7\%$ have osteoporosis.
    • Outcome: Osteoporosis.
    • Predictors: Vitamin D, calcium, weight, height, smoking, number of persons in household, age, US citizen status, education, gender. The following were considered nonlinear via rcs spline transformation with rms R-package [23] with 3 default knots:Vitamin D, calcium, age.
  • Smoking: Logistic regression model based on NHANES 2007/2008 [16] data using the nhanesA-package [2] with complete case analysis; $26.3\%$ do smoke.
    • Outcome: Smoking.
    • Predictors: Vitamin D, calcium, bmi, osteoporosis, number of persons in household, age, US citizen status, education, gender. The following were considered nonlinear via rcs spline transformation with rms R-package [23] with 3 default knots:Vitamin D, calcium, bmi, age.
E.2.2. Predictions $p_{i}$ entered in the BS
The predictions entered are functions of the $q_{i}$, and the following settings are considered:
  • Perfect: $p_{i}=q_{i}$, i.e., perfect predictions.
  • $+0.1$: $p_{i}=q_{i}+0.1$, i.e., slightly biased predictions.
  • $+Unif(-0.2,0.2)$: $p_{i}=q_{i}+X_{i}$, where $X_{i}\sim Unif(-0.2,0.2)$, i.e., fairly disturbed but unbiased predictions.
  • $+(1-2Bern(1/2))\cdot 0.1$: $p_{i}=q_{i}+(1-2X_{i})\cdot 0.1$, where $X_{i}\sim Bern(1/2)$, i.e., slightly disturbed but unbiased predictions.
In case the $p_{i}$ are smaller than zero then $p_{i}$ is set to zero, and if they are bigger than one set to one.
E.2.3. Sample Size
Sample sizes considered are $n=\{300,1000\}$.
E.2.4. Number of DGM Scenarios and Simulation Runs
[…]
  • $N=5000$ simulation repetitions per scenario.
E.3. Estimand/Target of Analysis
  • Mean of BS.
  • 2.5% and 97.5% quantile of BS.
E.4. Methods
E.4.1. Basis of Simulations
The simulation is run as a Monte-Carlo simulation, where for the two settings based on NHANES data, the $q_{i}$ are based on the predicted value of the logistic regression for the corresponding observation, and the $q_{i}$ are subsampled without replacement.
E.5. Performance Measures
  • Median of Brier score calculated as sample median.
  • 2.5% and 97.5% quantile of Brier score calculated as corresponding quantiles.
  • CIL and its comparison to the BS.
  • estimate the probability that $\bar{y}-\bar{y}^{2}>BS_{perf}$.
  • estimate the median of $\bar{y}-\bar{y}^{2}-BS_{perf}$.